# Topological Symmetry Breaking in Antagonistic Dynamics

### v2

@@
Fig. 3(a) and (b) show the temperature phase diagram of the system as a function of the fraction of negative links, $p$, for S and X defects on 2D square lattices, respectively. In particular, S defects correspond to the well-known Random Bond Ising Model (RBIM) or ±$\pm J$-model 47, 48 (see Methods). Results are achieved through extensive Monte Carlo simulations of spin-spin dynamics, running Eq. (3) on top of the signed structure. The inspection of the phase diagram confirms the well-known collective dynamical regimes for S defects: ferromagnetic and spin glass phases. Now, we characterize the eigenstates of the system using the corresponding signed eigenvectors of $\bar{L}$ to define topologically-based up and down spin islands. Then, we calculate the energy of the system at $T=0$ through the EA Hamiltonian as defined in Eq. (3). This confirms the degeneracy of energetic states at the critical fraction of negative links, $p_{\rm c}$, for both S and X defects. Fig. 3(a) and (b) report how the ground state energy gap disappears, leading to the degeneracy and, finally, inversion of the energetic levels of the system. Note that, for X defects, the diagram shows, however, distinct bifurcation lines separating new phases that are still unknown to the best of our knowledge. In particular, for low temperatures, increasing the fraction $p$, a first transition separates the ferromagnetic phase from a metastable one, where the system can get trapped into local minima up to the Ising critical temperature, $T_{\rm c}$ (see Fig. 3(c)). A second transition at higher $p$ leads to the spin glass phase, where the resilient and locally stable spin islands become unstable when thermal noise is applied. Notably, this new intermediate phase between the ferromagnetic and spin glass phases, the MM phase, exhibits multiple coexisting local islands of mesoscopic magnetic order. Note that, unlike the chaotic attractor “surfing” seen in spin glass phases at low temperatures, the MM phase allows access to highly diverse noise-resilient functional configurations.
@@
Note that, for S defects, the usually reported intriguing re-entrance of the ferroelectric phase to slightly higher values of $p$ when the thermal noise increases 49 can be ascribed to the breaking of small frustrated clusters due to noise effects. Specifically, thermal noise induces random flips with a single-spin interaction radius that is explicitly smaller than the size of the clustered defects. This effect first breaks small spin-clusters, reducing domain wall effects and making more robust the giant ferroelectric cluster.
@@
Fig. 3(c) and (d) illustrate the stability of different eigenstates of $\bar{L}$ as the system is heated up to different temperatures. We emphasize that, for X defects, the system can be heated up to the Ising critical temperature, where it is reset by noise, and allow for choosing other possible local minima. The same occurs at a much lower temperature, $T_{\rm SG}(N)\ll T_{\rm c}^{\rm Ising}$, for S defects. In the latter case, typical of spin glass behavior, the minima are locally stable, but the system can escape under low noise levels, exploring the phase space (see SI). We hypothesize that the spin-glass tree of states 9 is expected to be a superposition of the $N$ eigenvectors of the SL, or pure states, that represent the original set of local minima.
@@
It comes natural to introduce topologically-based local order parameters as a function of $T$ that effectively capture the emergent mesoscopic order when spontaneous symmetry breaking forms magnetized spin islands (see Methods). Fig. 3(e) shows the magnetization of the giant cluster of $|\lambda_{2}\rangle$ as initial state, as a function of $p$ and $T$ for the case of X defects. Hence, Fig. 3(f) and (g) show the spin overlap $\langle\lambda_{i}\,|\,\mathopen{}\sigma_{\infty}\rangle$ between the initial state and the long-time dynamics when the system has thermalized (see Methods). In particular, for the case of X defects, this evidences how when the system is initialized at $|\lambda_{2}\rangle$, it undergoes a second-order phase transition with a critical temperature a little below the Ising critical one, $T_{\rm c}\simeq 2.27$. This provides direct evidence of the stability of the corresponding eigenstate or initial configuration of the system. By contrast, Fig. 3(g) shows the narrow temperature range in which the selected starting eigenstates for S defects lose their stability, causing the system to start surfing across other attraction basis, i.e., pure states, or combinations thereof.
I Topological defects in heterogeneous architectures
To establish whether the SL is able to predict a percolative phase transition in heterogeneous systems, we have analyzed different dilute models and network structures by introducing a random fraction of S defects. We highlight that this approach generalizes the detection of frustrated dynamical phases to complex networks and generic disordered systems. Specifically, the results for Erdős-Rényi (ER) and Small World (SW) networks are reported in Fig. 4(a) and (b). Additionally, we have analyzed a particularly complex case with significant physical relevance: the 3D diluted cubic lattice, which has originally inspired dilute Ising models 50. Now, a fraction $f$ of the existent links are removed, remaining below the edge-percolation threshold to ensure the system is connected. We then assign positive and negative weights to the remaining interactions, randomly drawn from a Gaussian distribution with mean $\mu$ and variance $\sigma$. As illustrated in Fig. 4(c), we also predict the emergence of the frustrated phase as a function of $\sigma$. Finally, Fig. 4(d) shows the case of an ER network with X defects reconstructed with the local Ising order parameter.
@@ I Topological defects in heterogeneous architectures
Figure 4: Heterogeneous architectures. Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$ for a: (a) ER network with $\langle\kappa\rangle=10$, (b) Small World network from a 2D lattice (rewiring probability $p_{\rm rew}=0.05$) and (c) a diluted 3D lattice ($p_{\rm dil}=0.25$) with normally weighted edges $p\sim\mathcal{N}(1,\sigma)$. Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, $p$. (d) $T-p$ Phase diagram of the ER network with X defects reconstructed with the local Ising order parameter $m^{\rm GC}_{\infty}$ measured on $\operatorname{sign}(|\lambda_{2}\rangle)$.
@@ II Discussion and conclusions
We have solved a decades-long controversy surrounding the nature of spin glasses. We have explicitly shown how non-ergodic phases originated by competing interactions, such as the emergence of the spin glass phase at zero temperature, lie in a pure topological symmetry-breaking mechanism that goes beyond the considered dynamics. This mechanism exhibits a clear algebraic signature in the emergence of degeneracy of the ground state of the SL operator. Topological defects generate natural barriers in the discrete lattice that lead to the fragmentation of the system in mesoscopic regions already at low values of $p$. Naturally, such a mechanism strongly depends on the discrete nature of the underlying lattice, leading to different percolation thresholds even if the spatial dimension of the system remains unchanged. Hence, the SL framework offers a natural approach for detecting spin glass phases in any homogeneous or heterogeneous structure. This has profound implications for considering future dynamical Renormalization Group (RG) analysis as it directly acts on the Green function of the system by adding only a new term correcting $\operatorname{\nabla^{2}}$, which hampers the topological properties of the embedding space on top of which any field-theoretical approach is based.
@@
The percolative analysis presented here also permits an immediate interpretation of the long-standing debate of the upper critical dimension of spin glass systems 9. That is, on determining the dimensionality $d_{u}$ above which the critical behavior of spin glasses is described by mean-field theory. Suppose one assumes that the dimension of the percolative giant cluster determines the effective dimension where the spin system operates. In that case, $d=6$ is the upper dimension for the isotropic percolation universality class, which presents a fractal dimension $d_{f}=4$, which is precisely the upper dimension for the Ising universality class. Hence, specific spatial dimensions $d=4$ and $d=5$ are known to exhibit a fractal cluster with $d_{f}\approx 3.04$ and $d_{f}\approx 3.52$, respectively 51, making invalid the mean-field solution for Ising dynamics when competing interactions are considered.
@@ IV Methods
The traditional definition of the Laplacian operator, represented by $L=D-A$, presents serious issues when dealing with signed networks as it is no longer semi-definite positive. This is due to the fact that diagonal terms may take null or even negative values and $D^{-1/2}$ and, therefore, $L_{\rm RW}$ is no longer defined. To tackle this delicate point, a recent proposal has mathematically introduced the signed version of the Laplacian operator. It is grounded in sound physical principles, being capable of accurately capturing the complex and intricate diffusion dynamics of any network 34. In particular, the SL of a signed architecture can be defined as,
@@
In particular, Eq. (6) can be reduced, when considering only one negative link, to the problem of point-like particles diffusing in a uniform medium and being absorbed by a single spherical sink 52 (see further details in SI), with $m=m(\bm{x},t)$ the magnetization field,
@@
Through the extension of Eq. (7) to many sinks 52 one can find the following Green function,
@@
Once we can consider a semidefinite operator for signed networks, $\bar{L}$, it is possible, in terms of the network propagator, $\hat{K}=e^{-\tau\bar{L}}$, to generalize the ensemble of accessible information diffusion states 31, 33, 32, 53, namely,
@@
where $\rho(\tau)$ is tantamount to the canonical density operator in statistical physics (or to the functional over fields configurations) 35, 54, and $Z(\tau)=\sum_{i}^{N}e^{-\tau\lambda_{i}}$, being $\lambda_{i}$ the set of system’ eigenvalues. It is possible to, therefore, define the network entropy 31 through the relation
@@
Eq. (12) describes the heat capacity of the network, which can detect transition points that correspond to the intrinsic diffusion scales characteristic of the network. When condition ${\mathrm{d}C}/{\mathrm{d}\tau}|_{\tau^{*}}=0$ is met, it defines $\tau^{*}$ and reveals pronounced peaks that reflect a significant deceleration in information diffusion. If $C$ also shows a well-defined plateau, it represents direct evidence of a scale-invariant architecture 33, 36, 38 (such as, for example, 2D lattices). In this case, a peak at short times, in turn, independent of system size, captures the ultraviolet cutoff, $\Lambda$ 37. This cutoff represents a minimum scale that prevents divergence in usual RG calculations 55 and indicates that the system exhibits some kind of translational invariance if the subsequent plateau is present 38.
@@
The archetypal spin glass model is defined as the celebrated EA model 48, with the corresponding Hamiltonian
@@
$m^{(i)}_{\infty}(t)=\frac{1}{n^{(i)}_{\infty}}\stackrel{{\scriptstyle[}}{{j}}=1]{n^{(i)}_{\infty}}{\sum}\sigma_{j}(t)$ (17)
@@
In order to tighten the relation between the fundamental state of the SL and the ground state of the Hamiltonian Eq. (3) with the random interaction matrix $J_{ij}$, it is convenient to introduce the field theoretical formulation of the Ising model given by the Hubbard-Stratonovich transformation 56 for which
@@
for $k=1,2,....,N$, where $A_{ij}=J_{ij}/J$ is the adjacency matrix of the network of interactions. The self-consistent solution of the above equation is
@@
where $k_{i}=\sum_{j}|A_{ij}|$ is the coordination number of site $i$ ($k_{i}=2d$ for a $d-$dimensional square lattice). As shown in 57, $\lambda_{0}\leq\min_{i=1}^{N}k_{i}$. This implies that
@@
for all connected networks, including regular lattices. Eq. (21) is formally the same of Eq. (19) with the fundamental difference that the solution of the former has to be found on the $N-$dimensional spherical surface of radius $\sqrt{N}$ while for the latter has to found on the vertices of the $N-$dimensional cube defined by $\phi_{i}=\pm 1$ for each $i=1,2,...,N$ (note however that in both cases $|\bm{\psi}|=|\bm{\phi}|=\sqrt{N}$. This implies that $\operatorname{sign}[\psi_{i}]$ can differ from $\operatorname{sign}[\phi_{i}]$ in some node in particular where there are large fluctuations in the values of $\psi$ between neighboring nodes. Anyway, being $\bm{\phi}$ and $\bm{\psi}$ characterized by the same normalization condition, we expect that the binarized vector $\bm{\psi}_{\rm bin}$ of components $\psi_{i}/|\psi_{i}|$ shows a similar behavior with respect the binary vector $\bm{\phi}$ on a suitable coarse-grained scale as large as the frustration walls.

### v3

@@
Figure 3(a) and (b) show the temperature phase diagram of the system as a function of the fraction of negative links, $p$, for S and X defects on 2D square lattices, respectively. In particular, S defects correspond to the well-known Random Bond Ising Model (RBIM) or $\pm J$-model 48, 49 (see Methods). Results are achieved through extensive Monte Carlo simulations of spin-spin dynamics, running Eq. (3) on top of the signed structure. The inspection of the phase diagram confirms the well-known collective dynamical regimes for S defects: ferromagnetic and spin glass phases. Now, we characterize the eigenstates of the system using the corresponding signed eigenvectors of $\bar{L}$ to define topologically-based up and down spin islands. Then, we calculate the energy of the system at $T=0$ through the EA Hamiltonian as defined in Eq. (3). This confirms the degeneracy of energetic states at the critical fraction of negative links, $p_{\rm c}$, for both S and X defects. Fig. 3(a) and (b) report how the ground state energy gap disappears, leading to the degeneracy and, finally, inversion of the energetic levels of the system. Note that, for X defects, the diagram shows, however, distinct bifurcation lines separating new phases that are still unknown to the best of our knowledge. In particular, for low temperatures, increasing the fraction $p$, a first transition separates the ferromagnetic phase from a metastable one, where the system can get trapped into local minima up to the Ising critical temperature, $T_{\rm c}$ (see Fig. 3(c)). A second transition at higher $p$ leads to the spin glass phase, where the resilient and locally stable spin islands become unstable when thermal noise is applied. Notably, this new intermediate phase between the ferromagnetic and spin glass phases, the MM phase, exhibits multiple coexisting local islands of mesoscopic magnetic order. Note that, unlike the chaotic attractor “surfing” seen in spin glass phases at low temperatures, the MM phase allows access to highly diverse noise-resilient functional configurations.
@@
Note that, for S defects, the usually reported intriguing re-entrance of the ferroelectric phase to slightly higher values of $p$ when the thermal noise increases 50 can be ascribed to the breaking of small frustrated clusters due to noise effects. Specifically, thermal noise induces random flips with a single-spin interaction radius that is explicitly smaller than the size of the clustered defects. This effect first breaks small spin-clusters, reducing domain wall effects and making more robust the giant ferroelectric cluster.
@@
Figure 3(c) and (d) illustrate the stability of different eigenstates of $\bar{L}$ as the system is heated up to different temperatures. We emphasize that, for X defects, the system can be heated up to the Ising critical temperature, where it is reset by noise, and allow for choosing other possible local minima. The same occurs at a much lower temperature, $T_{\rm SG}(N)\ll T_{\rm c}^{\rm Ising}$, for S defects. In the latter case, typical of spin glass behavior, the minima are locally stable, but the system can escape under low noise levels, exploring the phase space (see SI). We hypothesize that the spin-glass tree of states 9 is expected to be a superposition of the $N$ eigenvectors of the SL, or or configurations that can approximate pure states, that represent the original set of local minima.
@@
It comes natural to introduce topologically-based local order parameters as a function of $T$ that effectively capture the emergent mesoscopic order when spontaneous symmetry breaking forms magnetized spin islands (see Methods). Fig. 3(e) shows the magnetization of the giant cluster of $|\lambda_{2}\rangle$ as initial state, as a function of $p$ and $T$ for the case of X defects. We examine a complementary mode to $|\lambda_{0}\rangle$ to probe spectral multiplicity and the presence of competing metastable structures. Hence, Fig. 3(f) and (g) show the spin overlap $\langle\lambda_{i}\,|\,\mathopen{}\sigma_{\infty}\rangle$ between the initial state and the long-time dynamics when the system has thermalized (see Methods). In particular, for the case of X defects, this evidences how when the system is initialized at $|\lambda_{2}\rangle$, it undergoes a second-order phase transition with a critical temperature a little below the Ising critical one, $T_{\rm c}\simeq 2.27$. This provides direct evidence of the stability of the corresponding eigenstate or initial configuration of the system. By contrast, Fig. 3(g) shows the narrow temperature range in which the selected starting eigenstates for S defects lose their stability, causing the system to start surfing across other attraction basis, i.e., pure states, or combinations thereof.
@@
Figure 4: Heterogeneous architectures. Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$ for a: (a) ER network with $\langle\kappa\rangle=10$, (b) Small World network from a 2D lattice (rewiring probability $p_{\rm rew}=0.05$) and (c) a diluted 3D lattice ($p_{\rm dil}=0.25$) with normally weighted edges $p\sim\mathcal{N}(1,\sigma)$. Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, pp. (d) TpT-p Phase diagram of the ER network with a fraction $p$ of X defects measured by the local Ising order parameter $m^{\rm GC}_{\infty}$, and using $\operatorname{sign}(|\lambda_{2}\rangle)$.
Figure 5: (a) Spectral gap ($\Delta$) and Energy Gap ($E_{0}-E_{1}$) versus $J_{0}/J$ for a fully connected network with symmetric Gaussian interactions, highlighting the closure of the gap at the transition point. (b) Activity decay for the antagonistic contact process with a fraction of inhibitory links, $p=0.15$. Note the slow decay region indicative of Griffiths phases that emerge beyond $p>0.1$. This confirms that the previously identified topological phase transition is consistent with the interpretation that it governs the divergence of dynamical timescales. This dynamical slowing down is analytically linked to the proliferation of vanishingly small eigenvalues in the Signed Laplacian spectrum, which emerge as topological defects disrupt the spectral gap. Thus, the topological fragmentation of the network provides a structural encoding of the broad distribution of relaxation timescales observed in the dynamics. Inset: Temporal decay for a 2D lattice with no inhibitory links ($p=0$), showing standard bistable behavior. Parameters: L=96.
Topological defects in heterogeneous architectures
To establish whether the SL is able to predict a percolative phase transition in heterogeneous systems, we have analyzed different dilute models and network structures by introducing a random fraction of S defects. We highlight that this approach generalizes the detection of frustrated dynamical phases to complex networks and generic disordered systems.
Specifically, the results for ER and Small World (SW) networks are reported in Fig. 4(a) and (b). Additionally, we have analyzed a particularly complex case with significant physical relevance: the 3D diluted cubic lattice, which has originally inspired dilute Ising models 51. Now, a fraction $f$ of the existent links are removed, remaining below the edge-percolation threshold to ensure the system is connected. We then assign positive and negative weights to the remaining interactions, randomly drawn from a Gaussian distribution with mean $\mu$ and variance $\sigma$. As illustrated in Fig. 4(c), we also predict the emergence of the frustrated phase as a function of $\sigma$. Finally, Fig. 4(d) shows the local Ising order parameter for the case of including a fraction $p$ of X defects in an ER network.
I Implication for E/I models of neural networks
Antagonistic interactions, central to the balance between excitation and inhibition in neural circuits, have long been a cornerstone of theoretical neuroscience. Moving beyond classical spin-lattice approaches that typically give rise to short-range spin glass models, we consider alternative formulations that provide a basis for more refined descriptions of neural computation. The first is the Sompolinsky–Crisanti–Sommers (SCS) model 52, which describes the dynamics of large recurrent networks with mixed excitatory and inhibitory synaptic interactions. This framework offers a minimal theoretical explanation for the intrinsically irregular, yet collectively organized population dynamics observed in cortical circuits 53. The model can be formulated in terms of continuous activity variables $x_{i}(t)$ that interact via quenched couplings $J_{ij}$ and are transformed through a nonlinear transfer function $\phi(\cdot)$; a complete treatment is provided in the Methods. Treating the connectivity as quenched disorder and applying dynamical mean-field theory 54, the model predicts a transition from a stable fixed-point regime to high-dimensional chaotic activity when the interaction strength exceeds a critical threshold. While the fully asymmetric case yields high-dimensional chaos, the onset of this instability is universally marked by the destabilization of the trivial fixed point (determined by the symmetric part of the interaction spectrum). As reported in Fig. 55(a), we observe that this critical point coincides with the closure of the spectral gap $\Delta$ of the underlying Signed Laplacian. Furthermore, the energy gap $E_{0}-E_{1}$ between the ground state and the first excited state of the effective energy landscape vanishes at the same threshold. This supports the view that the loss of stability in continuous rate models—regardless of whether the resulting phase is chaotic or glassy—is rooted in the same topological symmetry-breaking mechanism identified in discrete spin systems.
The second model we consider is a simple two-state (binary) neural network model 55, 56, in which each unit is either active ($s_{i}(t)=1$) or inactive ($s_{i}(t)=0$) at time $t$. The state of unit $i$ evolves according to the weighted activity of its $k$ neighbors, $\mathcal{P}_{i}=f\!\left(\Lambda_{i}=\frac{\gamma}{k}\sum_{j}\omega_{ij}s_{j}(t)\right)$, where $\omega_{ij}=\pm 1$ encodes excitatory and inhibitory interactions and $\gamma$ controls the interaction strength. To connect with the symmetric interaction framework used in this work, we consider a variant in which excitation and inhibition are assigned to links rather than nodes, so that $\omega_{ij}=\omega_{ji}$ and a fraction $\alpha$ of links are inhibitory. This link-based formulation preserves the global E/I balance while ensuring a symmetric coupling matrix, allowing us to isolate the collective effects of antagonistic interactions; full details are provided in the Methods section. Fig. 55(b) shows the temporal decay of the network activity from a fully active state. For low fractions of inhibition ($p<0.1$), the system exhibits a single critical point with rapid relaxation or simple bistability (see Inset). However, beyond the critical threshold $p\approx 0.1$, we observe the emergence of ’Griffiths phases’–regions characterized by an anomalously slow, power-law decay of activity. This phenomenon is not merely a dynamical artifact but a direct consequence of the spectral properties of the Signed Laplacian: as topological defects percolate, they induce a proliferation of degenerate eigenvalues. Since relaxation timescales are inversely proportional to these eigenvalues ($\tau\sim 1/\lambda$), the topological fragmentation of the network structurally guarantees the emergence of the diverse, long-lasting timescales often associated with criticality in neural circuits.
@@ II Discussion and conclusions
We provide a new perspective on a decades-long controversy surrounding the nature of spin glasses. We have explicitly shown how non-ergodic phases originated by competing interactions, such as the emergence of the spin glass phase at zero temperature, lie in a pure topological symmetry-breaking mechanism that goes beyond the considered dynamics. This mechanism exhibits a clear algebraic signature in the emergence of degeneracy of the ground state of the SL operator. Topological defects generate natural barriers in the discrete lattice that lead to the fragmentation of the system in mesoscopic regions already at low values of $p$. Naturally, such a mechanism strongly depends on the discrete nature of the underlying lattice, leading to different percolation thresholds even if the spatial dimension of the system remains unchanged. While the Signed Laplacian does not fully determine nonlinear dynamics, it provides a minimal structural backbone that constrains the landscape on which these dynamics unfold. Hence, the SL framework offers a natural approach for detecting spin glass phases in any homogeneous or heterogeneous structure. This has profound implications for considering future dynamical Renormalization Group (RG) analysis as it directly acts on the Green function of the system by adding only a new term correcting $\operatorname{\nabla^{2}}$, which hampers the topological properties of the embedding space on top of which any field-theoretical approach is based.
@@
The percolative analysis presented here also permits an immediate interpretation of the long-standing debate of the upper critical dimension of spin glass systems 9. That is, on determining the dimensionality $d_{u}$ above which the critical behavior of spin glasses is described by mean-field theory. Suppose one assumes that the dimension of the percolative giant cluster determines the effective dimension where the spin system operates. In that case, $d=6$ is the upper dimension for the isotropic percolation universality class, which presents a fractal dimension $d_{f}=4$, which is precisely the upper dimension for the Ising universality class. Hence, specific spatial dimensions $d=4$ and $d=5$ are known to exhibit a fractal cluster with $d_{f}\approx 3.04$ and $d_{f}\approx 3.52$, respectively 57, making invalid the mean-field solution for Ising dynamics when competing interactions are considered.
Percolation phenomena in correlated and higher-order networked systems have also been investigated in other contexts, such as multiplex and hypergraph models or reaction networks 58, 59. While these studies rely on different mechanisms, they are consistent with the broader view that structural organization can influence collective transitions and may motivate future extensions of the present framework.
One might ask whether other spectral metrics, such as the standard Laplacian or measures of diffusion capacity, could capture these phenomena. Standard graph operators, however, typically treat links as purely conductive channels (positive weights), thereby ignoring the crucial algebraic structure of antagonistic interactions. While metrics like the Diffusion Capacity describe transport efficiency, they fail to detect the frustration arising from negative cycles. The Signed Laplacian is unique in this regard: it naturally incorporates the sign of the interaction into the diffusion process, ensuring that the spectral gap $\Delta$ directly measures the structural balance of the network. Consequently, the Signed Laplacian provides a direct and physically interpretable link between the algebraic structure of antagonistic interactions and the effective energy landscape governing relaxation dynamics.
@@ IV Methods
The traditional definition of the Laplacian operator, represented by L=DAL=D-A, presents serious issues when dealing with signed networks as it is no longer semi-definite positive. This is due to the fact that diagonal terms may take null or even negative values and $D^{-1/2}$ and, therefore, $L_{\rm RW}$ is no longer defined. To tackle this delicate point, a recent proposal has mathematically introduced the signed version of the Laplacian operator. It is grounded in sound physical principles, being capable of accurately capturing the complex and intricate diffusion dynamics of any network 34. In particular, the SL of a signed architecture can be defined as,
@@
In particular, Eq. (6) can be reduced, when considering only one negative link, to the problem of point-like particles diffusing in a uniform medium and being absorbed by a single spherical sink 60 (see further details in SI), with $m=m(\bm{x},t)$ the magnetization field,
@@
Through the extension of Eq. (7) to many sinks 60 one can find the following Green function,
@@
Once we can consider a semidefinite operator for signed networks, $\bar{L}$, it is possible, in terms of the network propagator, $\hat{K}=e^{-\tau\bar{L}}$, to generalize the ensemble of accessible information diffusion states 31, 33, 32, 61, namely,
@@
where $\rho(\tau)$ is tantamount to the canonical density operator in statistical physics (or to the functional over fields configurations) 35, 62, and $Z(\tau)=\sum_{i}^{N}e^{-\tau\lambda_{i}}$, being $\lambda_{i}$ the set of system’s eigenvalues. It is possible to, therefore, define the network entropy 31 through the relation
@@
Eq. (12) describes the heat capacity of the network, which can detect transition points that correspond to the intrinsic diffusion scales characteristic of the network. When condition ${\mathrm{d}C}/{\mathrm{d}\tau}|_{\tau^{*}}=0$ is met, it defines $\tau^{*}$ and reveals pronounced peaks that reflect a significant deceleration in information diffusion. If $C$ also shows a well-defined plateau, it represents direct evidence of a scale-invariant architecture 33, 36, 38 (such as, for example, 2D lattices). In this case, a peak at short times, in turn, independent of system size, captures the shortest resolvable interaction scale (by analogy with an ultraviolet cutoff), denoted $\Lambda$ in lattice settings 37. This cutoff represents a minimum scale that prevents divergence in usual RG calculations 63 and indicates that the system exhibits some kind of translational invariance if the subsequent plateau is present 38.
@@
The archetypal spin glass model is defined as the celebrated EA model 49, with the corresponding Hamiltonian
@@
$m^{(i)}_{\infty}(t)=\frac{1}{n^{(i)}_{\infty}}\mathrel{{\mathop{\sum}\limits_{j=1}^{n^{(i)}_{\infty}}}}\sigma_{j}(t)$ (17)
@@
In order to tighten the relation between the fundamental state of the SL and the ground state of the Hamiltonian Eq. (3) with the random interaction matrix $J_{ij}$, it is convenient to introduce the field theoretical formulation of the Ising model given by the Hubbard-Stratonovich transformation 64 for which
@@
for k=1,2,.,Nk=1,2,....,N, where $A_{ij}=J_{ij}/J$ is the adjacency matrix of the network of interactions. The self-consistent solution of the above equation is
@@
where $k_{i}=\sum_{j}|A_{ij}|$ is the coordination number of site $i$ ($k_{i}=2d$ for a $d$-dimensional square lattice). As shown in 65, $\lambda_{0}\leq\min_{i=1}^{N}k_{i}$. This implies that
@@
for all connected networks, including regular lattices. Eq. (21) is formally the same of Eq. (19) with the fundamental difference that the solution of the former has to be found on the NN-dimensional spherical surface of radius $\sqrt{N}$ while for the latter has to found on the vertices of the NN-dimensional cube defined by $\phi_{i}=\pm 1$ for each i=1,2,,Ni=1,2,...,N (note however that in both cases $|\bm{\psi}|=|\bm{\phi}|=\sqrt{N}$. This implies that $\operatorname{sign}[\psi_{i}]$ can differ from $\operatorname{sign}[\phi_{i}]$ in some node in particular where there are large fluctuations in the values of $\psi$ between neighboring nodes. Anyway, being $\bm{\phi}$ and $\bm{\psi}$ characterized by the same normalization condition, we expect that the binarized vector $\bm{\psi}_{\rm bin}$ of components $\psi_{i}/|\psi_{i}|$ shows a similar behavior with respect the binary vector $\bm{\phi}$ on a suitable coarse-grained scale as large as the frustration walls.
Sompolinsky–Crisanti–Sommers (SCS) rate model
We consider a recurrent network of $N$ units with continuous activity variables $x_{i}(t)$, generated from effective inputs $h_{i}(t)$ through a nonlinear transfer function,
$x_{i}(t)=\phi\!\left(g\,h_{i}(t)\right),$ (22)
where $g$ is a gain parameter and $\phi(x)=\tanh(x)$ unless stated otherwise. The inputs evolve according to first-order relaxation dynamics,
$\frac{\mathrm{d}h_{i}(t)}{\mathrm{d}t}=-h_{i}(t)+\sum_{j=1}^{N}J_{ij}\,x_{j}(t),$ (23)
where $J_{ij}$ are quenched synaptic couplings drawn as
$J_{ij}\sim\mathcal{N}\!\left(0,\,\frac{g^{2}}{N}\right),\qquad J_{ii}=0.$ (24)
In the limit $N\to\infty$, the recurrent input to each neuron becomes an effective Gaussian process, and dynamical mean-field theory (DMFT) yields a self-consistent equation for the autocorrelation $C_{x}(t,t^{\prime})=\langle x(t)x(t^{\prime})\rangle$. This theory predicts a transition at $g_{c}=1$: for $g<g_{c}$ the network converges to a stable fixed point, while for $g>g_{c}$ the dynamics is chaotic.
The structure of the connectivity matrix determines whether the dynamics is relaxational. For generic (asymmetric) $J$,  Eq. (23) does not derive from an energy function, and the dynamics may exhibit chaotic behavior. However, when the couplings are symmetric, $J_{ij}=J_{ji}$, the dynamics becomes a gradient descent in an energy landscape,
$\frac{\mathrm{d}h_{i}(t)}{\mathrm{d}t}=-\frac{\partial\mathcal{E}}{\partial h_{i}}\Big|_{h(t)},$ (25)
with energy
$\mathcal{E}(h_{1},\dots,h_{N})$ $=\sum_{i=1}^{N}\int_{0}^{h_{i}}\mathrm{d}h\,\phi(gh)$ (26)
$\quad-\frac{1}{2}\sum_{i,j=1}^{N}J_{ij}\,\phi(gh_{i})\,\phi(gh_{j}).$
The first term encodes the single-unit activation nonlinearity, while the second term represents pairwise synaptic interactions.
Binary excitatory–inhibitory model with symmetric interactions
We also consider a binary-state network model in which each node is either active ($s_{i}(t)=1$) or inactive ($s_{i}(t)=0$) at discrete time $t$. Each node interacts with its $k$ neighbors through weighted inputs. The probability of activation is given by
$\mathcal{P}\big(s_{i}(t{+}1)=1\big)=f\!\left(\Lambda_{i}(t)\right),\qquad\Lambda_{i}(t)=\frac{\gamma}{k}\sum_{j}\omega_{ij}s_{j}(t),$
where $\gamma$ controls the global coupling strength and $\omega_{ij}\in\{\pm 1\}$ encodes excitatory (+1+1) or inhibitory (1-1) interactions. The activation function is taken to be piecewise linear:
$\begin{cases}0,&\Lambda<0,\\ \Lambda,&0\leq\Lambda\leq 1,\\[3.0pt] 1,&\Lambda>1.\end{cases}$
To connect directly with the symmetric interaction framework used in the present work, we consider a link-symmetrized variant in which inhibition and excitation are assigned to links rather than nodes. The network is taken to be random-regular and undirected, so that each node has exactly $k$ neighbors and $\omega_{ij}=\omega_{ji}\in\{\pm 1\}$. A fraction $p$ of links are inhibitory and (1p)(1{-}p) are excitatory, ensuring that the coupling matrix $(\omega_{ij})$ is symmetric by construction. The update rule remains unchanged, but inhibition is now encoded at the level of pairwise interactions rather than neuronal types, providing a minimal symmetric framework for a contact–process–like antagonistic coupling.